OpenStudy (anonymous):

find the square of the radical expression: the square root of:3x^9+3

OpenStudy (anonymous):

\[\sqrt{3x^9+3}\]

OpenStudy (anonymous):

\[(\sqrt{3x^9 +3})^2 = 3x^9+3\]

OpenStudy (anonymous):

In a more technically precise sense\[(\sqrt{3x^9+3})^2=|3x^9+3|\]

OpenStudy (anonymous):

The distinction is important if x<-1.